What is the pH of a sodium hydroxide solution at 25 oC in which the hydroxide ion concentration is 0.0026 mol/L? Is the solution acidic or basic?
1
Expert's answer
2019-08-19T04:52:02-0400

According to the pH concept:

pOH = -log[OH-]


As [OH-] = 0.0026 mol/L:

pOH = -log[0.0026] = 2.585


As pH + pOH = 14:

pH = 14 - pOH = 14 - 2.585 = 11.415 - the solution is basic as its pH is >7.


Answer: 11.415